Assume TC = 5Q2+ 200Q+ 110 is the cost function of a firm. This firm sells 30Q for a unit price of Rs. 10. What is the profit maximizing quantity of this firm?
1
Expert's answer
2020-06-09T17:39:41-0400

Here, In the question cost function is given as


"TC= 5 Q^2 + 200 Q+110"


here from above function we can say that


TFC=110


ATC (Q)="5Q+200+\frac{110}{Q}"


and TVC(Q)= "5Q^2+200Q"


AVC(Q)=200 +5Q


MC(Q) = 10Q+200

now we can say that if "P<200" , production is zero

here,

firms selling= 300

so profit

P= 10 Q+200

"Q= \frac{P-200}{10}" , this is profit maximization quantity of this firm